# On the spectral norm of Rademacher matrices

### v1

@@ 1. Introduction and main results
One of the basic issues of the random matrix theory are bounds on the spectral norm (largest singular value) of various families of random matrices. This question is very well understood for classical ensembles of random matrices [2], when one may use methods based on the large degree of symetry. Recently, a substantial progress was attained in the understanding of unhomogenous models [13], especially in the Gaussian case [9, 3]. However, there are still many open questions in this area, the one concerning Rademacher matrices is discussed here.
@@
where $(a_{i,j})$ is a deterministic matrix and $(\varepsilon_{i,j})_{i,j\geq 1}$ is the double indexed sequence of i.i.d. symmetric ±$\pm 1$ r.v’s.
@@
In [8, Theorem 1.1] it was shown that for any matrix $(a_{ij})$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|$ $\gtrsim\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$ (1)
@@
It was also conjectured that bound (1) may be reversed, i.e., for any scalar matrix $(a_{i,j})_{i,j\leq n}$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|$ $\lesssim\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$ (2)
The proof of [8, Remark 4.5], based on the permutation method from [9], shows that in order to establish (2) it is enough to show that for any submatrix $(b_{i,j})_{i,j\leq m}$ of $(a_{i,j})_{i,j\leq n}$ one has
${\mathbb{E}}\|(b_{i,j}\varepsilon_{i,j})_{i,j\leq m}\|\lesssim\max_{1\leq i\leq m}\|(b_{i,j})_{j}\|_{2}+\max_{1\leq j\leq m}\|(b_{i,j})_{i}\|_{2}+R_{B}(\mathrm{Log}\,m),$ (3)
@@
Theorem 1.
Inequality (3) holds if $b_{i,j}\in\{0,1\}$ for any $i,j$. As a consequence, for any $E\subset[n]\times[n]$,
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}\mathbbm{1}_{E}(i,j)\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$ (4)
Inequality (3) for $\{0,1\}$-weights is a consequence of the more general Theorem 5 below, applied to the symmetric $2m\times 2m$0,1}-matrix $A=\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$. Estimate (2) follows from (3) as in the proof of [8, Remark 4.5].
Remark 2.
[8, Proposition 1.4] gives an equivalent (up to a constant) form of $R_{A}(p)$ for $\{0,1\}$-matrices:
@@
Hence the first part of Theorem 1 gives a positive answer to the question posed by Ramon van Handel (private communication):
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\lesssim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
@@
One may also state the two-sided estimate (2) in the equivalent way as
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
@@
Remark 3.
Two-sided bound on moments of norms of Rademacher vectors [7] gives that for every $p\geq 1$,
@@
Thus, the first part of Theorem 1 might be equivalently stated as
$({\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|^{2\lfloor\mathrm{Log}\ n\rfloor})^{1/2\lfloor\mathrm{Log}n\rfloor}$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\max_{F\subset E,|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|.$ (5)
It is quite tempting to show (3) for symmetric sets $E$ via a combinatorial method, since for $n\times n$ symmetric matrix $A$ and $k=\lfloor\mathrm{Log}\,n\rfloor$, $\|A\|\sim(\mathrm{tr}(A^{2k}))^{1/{2k}}$. Such an approach worked for Gaussian matrices [4], but we were not able to apply it in the Rademacher case.
Remark 4.
Signed adjacency matrices were studied in [6] in connection with 2-lifts of graphs. [6, Lemma 3.1] shows that to each signed adjacency matrix of a graph $G$ one may associate the 2-lift of $G$ with the set of eigenvalues being the union of the eigenvalues of $G$ and of the signed matrix. Hence Theorem 1 provides an average uniform bound for new eigenvalues of random 2-lifts.
To state results for general matrices we need to introduce some additional notation. We associate to a symmetric matrix $(a_{i,j})_{i,j\leq n}$ a graph $G_{A}=([n],E_{A})$, where $(i,j)\in E_{A}$ iff $i\neq j$ and $a_{i,j}\neq 0$. By $d_{A}$ we denote the maximal degree of vertices in $G_{A}$. Observe that in the case of $\{0,1\}$-matrices $\sqrt{d_{A}}\|(a_{i,j})\|_{\infty}=\sqrt{d_{A}}\leq\max_{i}\|(a_{i,j})_{j}\|_{2}$.
Theorem 5.
@@
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}n)+d_{A}^{19/40}\|(a_{i,j})\|_{\infty}.$ (6)
Remark 6.
@@
${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{i\neq j\}}\varepsilon_{i,j})_{i,j}\|\leq d_{A}\|(a_{i,j})\|_{\infty}$ (7)
@@
The proof of Theorem 5 takes the most part of the paper. Bernoulli conjecture, formulated by Talagrand and proven in [5], states that to estimate a supremum of Bernoulli process one needs to decompose the index set into two parts and estimate supremum over the first part using the uniform bound and over the second part by the supremum of the Gaussian process. Unfortunately, there is no algorithmic method for making such a decomposition – a rule of thumb is that the uniform bound works well for large coefficients and Gaussian bound for small ones. We try to follow this informal recipe, decompose vectors $s,t\in B_{2}^{n}$ into almost "flat" parts and use the uniform bound when infinity norms of these parts differs. When they are of the same order we make some further technical adjustments (using properties of the graph $G_{A}$) and apply the Gaussian bound. The crucial tool in the estimating corresponding Gaussian process is an improvement of van Handel’s bound [12].
We postpone the details of the proof till the end of the paper and discuss now some of consequences of Theorem 5.
Theorem 7.
@@
Moreover, using again Theorem 5
@@
Remark 8.
In Theorems 5 and 7 we do not assume the symmetry of $(\varepsilon_{i,j})_{i,j}$. However analogous bounds holds for ${\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j})_{i,j}\|$, where $(\tilde{\varepsilon}_{i,j})_{i,j}$ is the symmetric Rademacher matrix (i.e., $\tilde{\varepsilon}_{i,j}=\tilde{\varepsilon}_{j,i}=\varepsilon_{i,j}$ for $i\geq j$), since
Obviously, $d_{A}\leq n$, so Theorem 7 (together with the standard symmetrization argument) implies that bounds (3) and (2) hold up double logarithms of $n$. However, decomposing matrix into two parts and using the Bandeira-van Handel bound one may derive conjectured upper bounds up to triple logarithms.
Theorem 9.
@@
Assume first that the matrix $(a_{i,j})$ is symmetric. Let $g_{i,j}$ be iid $\mathcal{N}(0,1)$ r.v’s. The result of Bandeira and van Handel [4] implies
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim{\mathbb{E}}\|(a_{i,j}g_{i,j})_{i,j\leq n}\|\lesssim\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}+\sqrt{\log n}\|(a_{i,j})\|_{\infty}.$ (8)
Put $M:=\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}$. Estimate (8) yields
@@
hence Theorem 7, applied to a matrix $(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|>M\mathrm{Log}^{-1/2}\,n\}})_{i,j\leq n}$ implies
@@
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\Bigl{(}\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n)\Bigr{)}.$ (9)
@@
Now, supppose that matrix $(a_{i,j})$ is arbitrary. Applying (9) to the symmetric $2n\times 2n$ matrix $\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$ we get the first part of the assertion.
The second part follows from the first one as in the proof of [8, Remark 4.5]. ∎

### v2

@@ 1. Introduction and main results
One of the basic issues of the random matrix theory are bounds on the spectral norm (largest singular value) of various families of random matrices. This question is very well understood for classical ensembles of random matrices [3], when one may use methods based on the large degree of symmetry. Recently, a substantial progress was attained in the understanding of inhomogeneous models [18], especially in the Gaussian case [12, 4]. However, there are still many open questions in this area, the one concerning Rademacher matrices is discussed here.
@@
where $(a_{i,j})$ is a deterministic matrix and $(\varepsilon_{i,j})_{i,j\geq 1}$ is the double indexed sequence of i.i.d. symmetric $\pm 1$ r.v’s.
@@
Let $g_{i,j}$ be independent $\mathcal{N}(0,1)$ random variables. We have
(1.1) ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|\lesssim{\mathbb{E}}\|(a_{i,j}g_{i,j})\|\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+\max_{j}\|(a_{i,j})_{i}\|_{2}+\sqrt{\log n}\max_{i,j}|a_{i,j}|,$
where the last bound was showned by Bandeira and van Handel [5]. The bound on the Gaussian matrices is sharp for $\{0,1\}$-weights; however, even in this case, estimate (1.1) often yields suboptimal bounds (see Section 2 below).
In [11, Theorem 1.1] it was shown that for any matrix $(a_{ij})$,
(1.2) ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|$ $\gtrsim\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$
@@
It was also conjectured that bound (1.2) may be reversed, i.e., for any scalar matrix $(a_{i,j})_{i,j\leq n}$,
(1.3) ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|$ $\lesssim\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$
The proof of [11, Remark 4.5], based on the permutation method from [12], shows that in order to establish (1.3) it is enough to show that for any submatrix $(b_{i,j})_{i,j\leq m}$ of $(a_{i,j})_{i,j\leq n}$ one has
(1.4) ${\mathbb{E}}\|(b_{i,j}\varepsilon_{i,j})_{i,j\leq m}\|\lesssim\max_{1\leq i\leq m}\|(b_{i,j})_{j}\|_{2}+\max_{1\leq j\leq m}\|(b_{i,j})_{i}\|_{2}+R_{B}(\mathrm{Log}\,m),$
@@
Theorem 1.1.
Inequality (1.4) holds if $b_{i,j}\in\{0,1\}$ for any $i,j$. As a consequence, for any $E\subset[n]\times[n]$,
(1.5) ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}\mathbbm{1}_{E}(i,j)\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$
Inequality (1.4) for $\{0,1\}$-weights is a consequence of the more general Theorem 1.6 below, applied to the symmetric $2m\times 2m$ {0,1}-matrix $A=\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$. Estimate (1.5) follows from (1.4) as in the proof of [11, Remark 4.5].
Remark 1.2.
In our results we do not assume the symmetry of $(\varepsilon_{i,j})_{i,j}$ even if we consider symmetric matrices $(a_{i,j})$. However analogous upper bounds holds for ${\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j})_{i,j}\|$, where $(\tilde{\varepsilon}_{i,j})_{i,j}$ is the symmetric Rademacher matrix (i.e., $\tilde{\varepsilon}_{i,j}=\tilde{\varepsilon}_{j,i}=\varepsilon_{i,j}$ for $i\geq j$), since
Moreover, the lower bound (1.2) remains valid if we replace $\varepsilon_{i,j}$ with $\tilde{\varepsilon}_{i,j}$.
Quantity $R_{A}(p)$ involves random variables $\varepsilon_{i,j}$. It may be expressed in terms of $a_{i,j}$ using two-sided bounds for $L_{p}$-norms of Rademacher sums (derived in [10] on the base of tail bounds [15])
$\Bigg{\|}\sum_{k=1}^{n}a_{k}\varepsilon_{k}\Bigg{\|}_{p}\sim\sum_{k\leq p}a_{k}^{*}+\sqrt{p}\Bigg{(}\sum_{k>p}(a_{k}^{*})^{2}\Bigg{)}^{1/2}\sim\sup\Bigg{\{}\sum_{k=1}^{n}a_{k}b_{k}\colon\|b\|_{\infty}\leq 1,\|b\|_{2}\leq\sqrt{p}\Bigg{\}},$
where $(a_{k}^{*})_{k=1}^{n}$ denotes the nonincreasing rearrangement of $(|a_{k}|)_{k=1}^{n}$. It is still unclear how to apply the above bounds to get a simple two-sided estimate for $R_{A}(p)$. However, in [11, Proposition 1.4], such an estimate was established for $\{0,1\}$-matrices:
@@
Hence Theorem 1.1 implies the following corollary – its first part provides a positive answer to the question posed by Ramon van Handel (private communication).
Corollary 1.3.
For any $E\subset[n]\times[n]$,
(1.6) ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\lesssim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
@@
and
(1.7) ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
@@
Example 2.7 below shows that one cannot reverse estimate (1.6) and a more involved form (1.7) of the two-sided estimate is necessary.
Remark 1.4.
Two-sided bounds on moments of norms of Rademacher vectors [9] give that for every $p\geq 1$,
@@
Thus, estimate (1.6) might be equivalently stated as
(1.8) $({\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|^{2\lfloor\mathrm{Log}\ n\rfloor})^{1/2\lfloor\mathrm{Log}n\rfloor}$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\max_{F\subset E,|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|.$
It is quite tempting to show (1.8) for symmetric sets $E$ via a combinatorial method, since for $n\times n$ symmetric matrix $A$ and $k=\lfloor\mathrm{Log}\,n\rfloor$, $\|A\|\sim(\mathrm{tr}(A^{2k}))^{1/{2k}}$. Such an approach worked for Gaussian matrices [5], but we were not able to apply it in the Rademacher case.
Remark 1.5.
Signed adjacency matrices were studied in [7] in connection with 2-lifts of graphs. [7, Lemma 3.1] shows that to each signed adjacency matrix of a graph $G$ one may associate the 2-lift of $G$ with the set of eigenvalues being the union of the eigenvalues of $G$ and of the signed matrix. Hence Theorem 1.1 provides an average uniform bound on new eigenvalues of random 2-lifts.
To state results for general matrices we need to introduce some additional notation. We associate to a symmetric matrix $(a_{i,j})_{i,j\leq n}$ a graph $G_{A}=([n],E_{A})$, where $(i,j)\in E_{A}$ iff $i\neq j$ and $a_{i,j}\neq 0$. By $d_{A}$ we denote the maximal degree of vertices in $G_{A}$. Observe that in the case of $\{0,1\}$-matrices $\sqrt{d_{A}}\|(a_{i,j})\|_{\infty}=\sqrt{d_{A}}=\max_{i}\|(a_{i,j})_{j}\|_{2}$.
Theorem 1.6.
@@
(1.9) ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n)+d_{A}^{19/40}\|(a_{i,j})\|_{\infty}.$
Remark 1.7.
@@
(1.10) ${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{i\neq j\}}\varepsilon_{i,j})_{i,j}\|\leq d_{A}\|(a_{i,j})\|_{\infty}$
@@
The proof of Theorem 1.6 takes the most part of the paper. Here we briefly sketch the main ideas of this proof. Bernoulli conjecture, formulated by Talagrand and proven in [6], states that to estimate a supremum of the Bernoulli process one needs to decompose the index set into two parts and estimate supremum over the first part using the uniform bound and over the second part by the supremum of the Gaussian process. Unfortunately, there is no algorithmic method for making such a decomposition – a rule of thumb is that the uniform bound works well for large coefficients and the Gaussian bound for small ones. We try to follow this informal recipe, decompose vectors $s,t\in B_{2}^{n}$ into almost "flat" parts and use the uniform bound when infinity norms of these parts are far apart. When they are of the same order we make some further technical adjustments (using properties of the graph $G_{A}$) and apply the Gaussian bound. The crucial tool used to estimate the corresponding Gaussian process is an improvement of van Handel’s bound [17], provided in Section 2.1.
We postpone the details of the proof till the end of the paper and discuss now some consequences of Theorem 1.6.
Theorem 1.8.
@@
Moreover, using again Theorem 1.6
@@
Obviously, $d_{A}\leq n$, so Theorem 1.8 (together with the standard symmetrization argument) implies that bounds (1.4) and (1.3) hold up double logarithms of $n$. However, decomposing matrix into two parts and using the Bandeira-van Handel bound one may derive conjectured upper bounds up to triple logarithms.
Theorem 1.9.
@@
Assume first that the matrix $(a_{i,j})$ is symmetric. Put $M:=\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}$. Estimate (1.1) yields
@@
hence Theorem 1.8, applied to a matrix $(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|>M\mathrm{Log}^{-1/2}\,n\}})_{i,j\leq n}$ implies
@@
(1.11) ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\Bigl{(}\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n)\Bigr{)}.$
@@
Now, supppose that matrix $(a_{i,j})$ is arbitrary. Applying (1.11) to the symmetric $2n\times 2n$ matrix $\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$ we get the first part of the assertion.
The second part follows from the first one as in the proof of [11, Remark 4.5]. ∎